# The effect of matter discreteness on gravitational wave propagation in post-geometrical optics

### v1

@@ 2.1 Post-geometrical approximation
The tensor $M_{\alpha\beta\mu\nu}$ is separately symmetric in the first and the second set of indices, and $M_{\alpha\beta\mu\nu}=M_{\mu\nu\alpha\beta}$. Equation (2.4) has the correct limit around the Friedmann–Lemaître–Robertson–Walker metric, unlike putting $\delta T_{\alpha\beta}=0$, although the difference is only important for wavelengths close to the cosmological scale. It can be compared to the equation of motion for the vector potential $A^{\alpha}$ of the electromagnetic field in the Lorenz gauge in curved spacetime discussed in [8], $\Box A^{\alpha}-R^{\alpha}{}_{\beta}A^{\beta}=0$. In the gravitational wave case, the Weyl tensor enters in addition to the Ricci tensor, and unlike for electromagnetic waves, direct coupling to the Riemann tensor is present in all gauges.
@@
In the geometrical optics approximation the field has the local plane wave form [1], [2] (page 42)
@@
where the amplitude is expanded in a series of the constant parameter $\epsilon\ll 1$, which represents the ratio of the wavelength to other relevant scales that are considered large. As in [8], we consider a post-geometrical optics approximation where the spacetime curvature radius is not assumed to be large compared to the wavelength. (The curvature was also included in [12], where the Riemann tensor was taken to be a perturbative correction and only the vacuum case was considered.) Inserting the expansion (2.6) into the equation of motion (2.4) gives
@@
where $|\vec{k}|$ is the spatial wavenumber, and as usual the wavelength is $2\pi/|\vec{k}|$. (For a summary of similar dispersion relations for gravitational waves in matter found in various approximations, see [9]; [13] is closest to the present work.) For $E^{2}<\frac{2\rho}{3M_{{}_{\mathrm{Pl}}}^{2}}$, the wave cannot propagate inside matter, and the local plane wave approximation is not valid.
@@
where on the second line we have used (2.5), (2.14), and (2.16). The polarisation vector is not in general parallel transported along the light ray, but exhibits circular birefringence.
@@ 2.3 Discussion
The relative correction to the angular diameter distance $D_{A}$ from (2.24) is $\frac{5}{4608\pi^{13/2}}\frac{m^{6}}{M_{{}_{\mathrm{Pl}}}^{2}E^{4}}=6\times 10^{-7}\frac{m^{6}}{M_{{}_{\mathrm{Pl}}}^{2}E^{4}}$. (The equation for $D_{A}$ is (2.49) of [8], with a different numerical coefficient for the curvature term.) This becomes of order unity for wave energy $E_{D}=0.03m(m/M_{{}_{\mathrm{Pl}}})^{1/2}=0.02\text{ eV}(m/\text{GeV})^{3/2}$, or equivalently for particle mass $m=10(E/\text{eV})^{2/3}$ GeV $4(f/\text{Hz})^{2/3}$ eV, where $f$ is the frequency of the gravitational wave. The maximum curvature is of the order of the maximum graviton mass due to a particle, $E_{\text{C}}\equiv\sqrt{2\rho_{\text{max}}/(3M_{{}_{\mathrm{Pl}}}^{2})}=0.01m^{2}/M_{{}_{\mathrm{Pl}}}=5\times 10^{-12}(m/\text{GeV})^{2}$ eV. Equivalently, the particle mass for which the curvature scale is equal to the graviton energy is $m=30(f/\text{Hz})^{1/2}$ MeV. In general, wave optics effects are important when the wavelength $2\pi/E$ (assuming $E\gg E_{\text{C}}$) is larger than the Schwarzschild radius $r_{\text{s}}=m/(4\pi M_{{}_{\mathrm{Pl}}}^{2})$. The limiting energy is $E_{\text{WO}}=8\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}/m$. We denote $\varepsilon\equiv m/M_{{}_{\mathrm{Pl}}}$; as we consider particles such that classical gravity is a good approximation, $\varepsilon\ll 1$. So the hierarchy of scales is $E_{\text{WO}}=80m\varepsilon^{-2}\gg E_{D}=0.03m\varepsilon^{1/2}\gg E_{\text{C}}=0.01m\varepsilon$. The gravitational waves impacted by a single particle at a time are always deep in the wave optics regime, since the Schwarzschild radius goes down with decreasing mass. However, the small mass also means that the deflection angle is tiny, and wave effects are expected to be negligible. The graviton mass remains small, $\sim\varepsilon^{1/2}E$, even when the correction to the distance becomes significant, so the local plane wave approximation is not obviously wrong.
@@
Gravitational wave energy is $E=4\times 10^{-15}f$/Hz eV. For the local plane wave approximation to be valid, $E\gg E_{\text{C}}$ is necessary. The lowest LIGO–Virgo–KAGRA frequency $\sim 10$ Hz corresponds to $E=4\times 10^{-14}$ eV, so the approximation is reliable for $m\lesssim$ 100 MeV, and the distance correction is large for $m\gtrsim 10$ eV. For LISA, the smallest frequency is $\sim 10^{-4}$ Hz, so the corresponding mass range is 0.1 MeV $\gtrsim m\gtrsim 0.01$ eV. Pulsar timing arrays with the lowest frequency $\sim 10^{-9}$ Hz correspond to the mass range 1 keV $\gtrsim m\gtrsim 10^{-6}$ eV, but if the distances of stochastic sources are not determined, there is no measurable signal. In [8], dark matter masses $10^{4}$ GeV $\gtrsim m\gtrsim 100$ MeV were excluded by lack of observed CMB distortion.
@@
In contrast to [8], where only dark matter was relevant as light cannot propagate freely through charges, for gravitational waves also electrons and baryons have to be taken into account. Their correction to the distance is weighed by their contribution to the total matter density, which is $0.2$ for baryons and $10^{-5}$ for electrons. However, as the correction is proportional to $E^{-4}$, this has only a small effect on the limiting energy. Baryon masses are above the regime of validity for LIGO–Virgo–Kagra frequencies (and obviously for all smaller frequencies), so the post-geometrical optics approximation is clearly not valid, and no conclusions can be drawn. In contrast, electron mass is so small that the induced graviton mass remains negligible at the lowest frequency observed by LIGO–Virgo–KAGRA, so the local plane wave approximation is not obviously incorrect. The resulting correction to $D_{A}$ is overwhelmingly large. From (2.49) of [8] (replacing the numerical coefficient with the one derived here for gravitational waves), the equation of motion for the small redshifts $z\ll 1$ relevant for LIGO–Virgo–KAGRA is $\frac{\mathrm{d}^{2}D_{A}}{\mathrm{d}z^{2}}=-\beta^{2}D_{A}$, where $\beta^{2}\equiv\frac{5}{3072\pi^{13/2}}\Omega_{e0}\frac{m_{e}^{6}}{M_{{}_{\mathrm{Pl}}}^{2}E^{4}}=3\times 10^{22}$, where $\Omega_{e0}$ is the electron density parameter, $m_{e}$ is electron mass, and $E=4\times 10^{-14}$ eV corresponds to 10 Hz. This leads to caustic formation at distance $\pi\beta^{-1}H_{0}^{-1}\sim 0.1$ pc (where $H_{0}$ is the Hubble constant), well below any cosmological scales. The local plane wave approximation and hence the post-geometrical approximation is thus strongly violated. (The distance is so small that the approximations leading to (2.49) of [8] that assume the cosmic average mass density are not valid, but this makes little difference to the conclusion.) Such caustics could have a strong effect on the distance-redshift relation on the LIGO–Virgo–KAGRA observations, but work beyond the post-geometrical optics approximation is needed to check their existence and impact.

### v2

@@ 2.1 Post-geometrical approximation
The tensor $M_{\alpha\beta\mu\nu}$ is separately symmetric in the first and the second set of indices, and $M_{\alpha\beta\mu\nu}=M_{\mu\nu\alpha\beta}$. Equation (2.4) has the correct limit around the Friedmann–Lemaître–Robertson–Walker metric, unlike putting $\delta T_{\alpha\beta}=0$, although the difference is only important for wavelengths close to or larger than the Hubble length (for zero spatial curvature). It can be compared to the equation of motion for the vector potential $A^{\alpha}$ of the electromagnetic field in the Lorenz gauge in curved spacetime discussed in [8], $\Box A^{\alpha}-R^{\alpha}{}_{\beta}A^{\beta}=0$. In the gravitational wave case, the Weyl tensor enters in addition to the Ricci tensor, and unlike for electromagnetic waves, direct coupling to the Riemann tensor is present in all gauges.
@@
The geometrical optics approximation is based on the assumption that the field has the local plane wave form [1], [2] (page 42)
@@
where the amplitude is expanded in a series of the constant parameter $\epsilon\ll 1$, which represents the ratio of the wavelength to other relevant scales that are considered large. When the curvature radius is much larger than the wavelength, the leading terms give geometrical optics, and subleading terms give energy- and polarisation-dependent propagation, including the gravitational Hall effect for both light [12, 13, 14] and gravitational waves [15, 16, 17]. As in [8], we instead consider a post-geometrical optics approximation where the spacetime curvature radius is not assumed to be large compared to the wavelength. (The curvature was also included in [18], where the Riemann tensor was taken to be a perturbative correction and only the vacuum case was considered.) Inserting the expansion (2.6) into the equation of motion (2.4) gives
@@
where $|\vec{k}|$ is the spatial wavenumber, and as usual the wavelength is $2\pi/|\vec{k}|$. (For a summary of similar dispersion relations for gravitational waves in matter found in various approximations, see [9]; [19] is closest to the present work.) For $E^{2}<\frac{2\rho}{3M_{{}_{\mathrm{Pl}}}^{2}}$, the wave cannot propagate inside matter, and the local plane wave approximation is not valid. The situation is different from the case of massive gravitons considered in modified gravity: here the mass is non-zero only in the fraction of space where matter energy density is non-zero [20]. Thus the gradient of the mass is non-zero, leading to non-geodesic motion.
@@
where on the second line we have used (2.5), (2.14), and (2.16). The polarisation vector is not in general parallel transported along the light ray, but exhibits circular birefringence. It has been argued that the Weyl term generates extra polarisation modes even in vacuum when treated perturbatively as a higher-order term in $\epsilon$ [18, 21]. This is similar to how an electromagnetic field propagating through a medium can generate the scalar component $A^{0}$ even if it is initially set to zero. In neither case does the number of degrees of freedom grow: any new components are generated by the two propagating modes only. Understanding propagation beyond geometrical optics is important for observationally distinguishing between such possible novel effects in general relativity and modified gravity that can involve new physical degrees of freedom [22, 23].
@@ 2.3 Discussion
The relative correction to the angular diameter distance $D_{A}$ from (2.24) is $\frac{5}{4608\pi^{13/2}}\frac{m^{6}}{M_{{}_{\mathrm{Pl}}}^{2}E^{4}}=6\times 10^{-7}\frac{m^{6}}{M_{{}_{\mathrm{Pl}}}^{2}E^{4}}$. (The equation for $D_{A}$ is (2.49) of [8], with a different numerical coefficient for the curvature term.) This becomes of order unity for wave energy $E_{D}=0.03m(m/M_{{}_{\mathrm{Pl}}})^{1/2}=0.02\text{ eV}(m/\text{GeV})^{3/2}$, or equivalently for particle mass $m=10(E/\text{eV})^{2/3}$ GeV $4(f/\text{Hz})^{2/3}$ eV, where $f$ is the frequency of the gravitational wave. The maximum curvature is of the order of the maximum graviton mass due to a particle, $E_{\text{C}}\equiv\sqrt{2\rho_{\text{max}}/(3M_{{}_{\mathrm{Pl}}}^{2})}=0.01m^{2}/M_{{}_{\mathrm{Pl}}}=5\times 10^{-12}(m/\text{GeV})^{2}$ eV. Equivalently, the particle mass for which the curvature scale is equal to the graviton energy is $m=30(f/\text{Hz})^{1/2}$ MeV. In general, wave optics effects are important when the wavelength $2\pi/E$ (assuming $E\gg E_{\text{C}}$) is larger than the Schwarzschild radius $r_{\text{s}}=m/(4\pi M_{{}_{\mathrm{Pl}}}^{2})$. The limiting energy is $E_{\text{WO}}=8\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}/m$. We denote $\varepsilon\equiv m/M_{{}_{\mathrm{Pl}}}$; as we consider particles such that classical gravity is a good approximation, $\varepsilon\ll 1$. So the hierarchy of scales is $E_{\text{WO}}=80m\varepsilon^{-2}\gg E_{D}=0.03m\varepsilon^{1/2}\gg E_{\text{C}}=0.01m\varepsilon$. The gravitational waves impacted by a single particle at a time are always deep in the wave optics regime, since the Schwarzschild radius goes down with decreasing mass. However, the small mass also means that the deflection angle is tiny. The graviton mass remains small, $\sim\varepsilon^{1/2}E$, even when the correction to the distance becomes significant, so the local plane wave approximation is not obviously wrong.
@@
Gravitational wave energy is $E=4\times 10^{-15}f$/Hz eV. For the local plane wave approximation to be valid, $E\gg E_{\text{C}}$ is necessary. The lowest LIGO–Virgo–KAGRA frequency $\sim 10$ Hz corresponds to $E=4\times 10^{-14}$ eV, so the approximation is reliable for $m\lesssim$ 100 MeV, and the distance correction is large for $m\gtrsim 10$ eV. For LISA, the smallest frequency is $\sim 10^{-4}$ Hz, so the corresponding mass range is 0.1 MeV $\gtrsim m\gtrsim 0.01$ eV. Pulsar timing arrays with the lowest frequency $\sim 10^{-9}$ Hz correspond to the mass range 1 keV $\gtrsim m\gtrsim 10^{-6}$ eV, but there is a measurable signal only if distance to the sources can be determined. In [8], dark matter masses $10^{4}$ GeV $\gtrsim m\gtrsim 100$ MeV were excluded by lack of observed CMB distortion.
@@
In contrast to [8], where only dark matter was relevant as light cannot propagate freely through charges, for gravitational waves also electrons and baryons have to be taken into account. Their correction to the distance is weighed by their contribution to the total matter density, which is $0.2$ for baryons and $10^{-5}$ for electrons. However, as the correction is proportional to $E^{-4}$, this has only a small effect on the limiting energy. Baryon masses are above the regime of validity for LIGO–Virgo–Kagra frequencies (and obviously for all smaller frequencies), so the post-geometrical optics approximation is clearly not valid, and no conclusions can be drawn. In contrast, electron mass is so small that the induced graviton mass remains negligible at the lowest frequency observed by LIGO–Virgo–KAGRA, so the local plane wave approximation is not obviously incorrect. The resulting correction to $D_{A}$ is overwhelmingly large. From (2.49) of [8] (replacing the numerical coefficient with the one derived here for gravitational waves), the equation of motion for small redshifts $z\ll 1$ is $\frac{\mathrm{d}^{2}D_{A}}{\mathrm{d}z^{2}}=-\beta^{2}D_{A}$, where $\beta^{2}\equiv\frac{5}{3072\pi^{13/2}}\Omega_{e0}\frac{m_{e}^{6}}{M_{{}_{\mathrm{Pl}}}^{2}E^{4}}=3\times 10^{22}$, where $\Omega_{e0}$ is the electron density parameter, $m_{e}$ is electron mass, and $E=4\times 10^{-14}$ eV corresponds to 10 Hz. This leads to caustic formation at distance $\pi\beta^{-1}H_{0}^{-1}\sim 0.1$ pc (where $H_{0}$ is the Hubble constant), well below any cosmological scales. The local plane wave approximation and hence the post-geometrical approximation is thus strongly violated. (The distance is so small that the approximations leading to (2.49) of [8] that assume the cosmic average mass density are not valid, but this makes little difference to the conclusion.) Such caustics could have a strong effect on the distance-redshift relation on the LIGO–Virgo–KAGRA observations, but work beyond the post-geometrical optics approximation is needed to check their existence and impact.